OpenStudy (davidfedrick):

A conservationist determines that a particular beach is eroding at a rate of 1.1% each year. When writing an explicit formula to represent the amount of beach remaining each year, which value should she use as the common ratio? 0.011 0.089 0.989 1.011

OpenStudy (tkhunny):

Erosion Rate = 1.1% = 0.011 Stable Rate = 1 - Erosion Rate = 1 - 0.011 = ?? f(0) = 1 f(1) = 1 * Stable Rate f(2) = 1 * (Stable Rate)^2 Just build it. Where are you stuck? Show your work.